OpenStudy (anonymous):

Identify the solution of the equation 4|2x – 1| = 12.

myininaya (myininaya):

divide both sides by 4 |2x-1|=3 so we have 2x-1=3 2x-1=-3 2x=4 2x=-2 x=2 x=-1

OpenStudy (anonymous):

You rule! Thanks!

myininaya (myininaya):

np

OpenStudy (saifoo.khan):

SHe is correct.